Let (h,k) be a fixed point where h,k>0. a straight line passing through this point cuts the positive direction of coordinate axes at point P and Q. Find the minimum area of the triangle OPQ, O being the origin.


I FOUND THE AREA TO BE=1/2*[2hk+(h)squared*tan(theta)+(k)squared*cot(theta)]


but the answer is 2hk..... how the area can be minimised???


 

Grade:12

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Parth Tiwari:

Ans:- Let the slope is m. then the equation of the line is y-k=m(x-h) It cuts the co ordinate axix at the points

A={(k-mh), 0} and B={0,(h-m/k)} hence the area is Δ=1/2(k-mh)(h-m/k) then applying min condition for the variable m we get the min area is =2hk.

 

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Parth Tiwari !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur